# The systemic recoil velocity distribution and the scale height of isolated millisecond pulsar systems: Implications on neutron star retention fractions in star clusters

### v1

@@ 5 The indicative millisecond pulsar retention fractions in star clusters
[…]
@@ 6 Summary and Prospects
This work derives the systemic recoil velocity ($v_{\mathrm{sys}}$) distribution of isolated millisecond pulsar (MSP) systems from the observed $v_{\mathrm{l}}$ (the Galactic-longitude component of the transverse peculiar velocity), assuming that 1) $v_{\mathrm{l}}$ is statistically stable over time (the “stable-$v_{\mathrm{l}}$” assumption), and 2) the $\bm{v}_{\mathrm{sys}}$ directions are uniformly distributed (the “isotropic-$\bm{v}_{\mathrm{sys}}$” assumption). We found that the $v_{\mathrm{sys}}$ distribution can be approximated by a linear combination of three Maxwellian distributions. Using dynamical population synthesis (DPS), we validated the “stable-$v_{\mathrm{l}}$” and evaluated the initial and current Galaxy-wide scale heights of isolated MSP systems by matching the observed Galactic heights and $v_{\mathrm{l}}$ (the Galactic-latitude component of the transverse peculiar velocity) to the DPS counterparts. The MSP $v_{\mathrm{sys}}$ distribution reported in this work predicts relatively large MSP retention fractions that may account for the large number of MSPs retained in globular clusters.
@@ Acknowledgements.
HD appreciates the EACOA Fellowship awarded by the East Asia Core Observatories Association.
@@ Appendix A Selection of the millisecond pulsar sample
We selected the MSP sample from the PSRCAT Catalogue1 (Manchester et al. 2005) with the help of the psrqpy Python package (Pitkin 2018). The MSPs were searched using the following criteria:
@@
    $P_{0}<\leavevmode\nobreak$60 ms, where $P_{0}$ represents the spin periods of MSPs, and
@@
    $\dot{P}_{0}<2\times 10^{-17}\,{\mathrm{s\leavevmode\nobreak\ s^{-1}}}$, where $\dot{P}_{0}$ denotes the time derivative of $P_{0}$, and
@@
    both proper motion magnitude and parallax are of $>3\,\sigma$ significance, with the uncertainty in the proper motion magnitude calculated through the propagation of uncertainties in the right ascension (RA) and the declination components.
@@
There is no clear $P_{0}$ boundary between MSPs and other pulsars. Various $P_{0}$ upper limits up to $\sim 100$ ms have been adopted in different studies of MSPs (e.g. Manchester 2017; Spiewak et al. 2022; Ding et al. 2023). In this work, we adopted the indicative $P_{0}$ upper limit of 60 ms. Additionally, we followed Spiewak et al. (2022), applying the filter $\dot{P}_{0}<2\times 10^{-17}\,\mathrm{s\leavevmode\nobreak\ s^{-1}}$ to exclude recently born ($\lesssim 10$ kyr old) pulsars (such as the Crab pulsar with 34 ms spin period), except for PSR J1417--4402 that does not have a reported $\dot{P}_{0}$. Despite the $\dot{P}_{0}$ being unknown to us, PSR J1417--4402 is suggested to have entered the late recycling phase (Camilo et al. 2016), therefore likely being kinematically similar to many other MSPs.
@@
As mentioned in Section 2, we look for isolated MSP systems with $>3\,\sigma$ parallax determinations. However, the Gaia-based astrometric results registered in the PSRCAT Catalogue come from Jennings et al. (2018) based on the Gaia Data Release (DR) 2 (instead of the latest Gaia DR3). Therefore, in practice, we conducted the sample search in three steps. In the first step, we searched for MSPs with a discounted parallax significance threshold of $3\times 0.8=2.4$, where 0.8 is the expected ratio of parallax precision between the Gaia DR3 and the Gaia DR2. In the second step, we updated the astrometric results of Jennings et al. (2018) to Gaia DR3 utilizing the astroquery package (Ginsburg et al. 2019). In the final step, we removed MSPs with insignificant ($<3\,\sigma$) determinations of parallaxes or proper motions. Compared to parallaxes, proper motions are usually determined at higher significance. The proper motion filter (as part of the criterion iv) only removes one MSP (i.e., PSR J1811--2405) that is too close to the ecliptic plane to acquire precise proper motion in the declination direction with pulsar timing (Ng et al. 2020). Through the three-step procedure, we reached a sample of 65 isolated MSP systems, which are compiled in Table 2 along with their most precise proper motion and parallax determinations.
@@
After calculating the transverse peculiar velocities $v_{\perp}$ of the 65 MSPs (see Section 2), we found no correlation between $P_{0}$ and $v_{\perp}$ (with Pearson coefficient of $-0.06$), hence indicating negligible selection effect of $P_{0}$ on $v_{\perp}$.
@@
Table 2: Distances $d$, Galactic heights $z$, transverse peculiar velocities $\textit{{v}}_{\perp}\equiv\left(v_{\mathrm{l}},v_{\mathrm{b}}\right)$, and their magnitudes $v_{\perp}$ of the 65 selected millisecond pulsars calculated from their published proper motions ($\mu_{\alpha}$ and $\mu_{\delta}$) and parallaxes $\varpi$, where 1) negative $z$ values correspond to negative Galactic latitudes $b$, and 2) $v_{\mathrm{b}}$ and $v_{\mathrm{l}}$ represent the transverse peculiar velocity components along the directions of $b$ and Galactic longitudes $l$, respectively.

### v3

@@ 5 The indicative millisecond pulsar retention fractions in star clusters
At the time of writing, about 345 pulsars — predominantly MSPs, have been found to be retained in 45 globular clusters (see the “Pulsars in globular clusters” Catalogue [^4] https://www3.mpifr-bonn.mpg.de/staff/pfreire/GCpsr.html). The high NS retention fraction of over 10% to 20% in globular clusters has been difficult to reconcile with the relatively large NS peculiar velocities (Cordes97; Toscano99; Hobbs05; Gonzalez11) that are believed to easily exceed the generally small globular cluster central escape velocities ($\sim 50$$\mathrm{km\penalty 10000\ s^{-1}}$); this NS retention problem in globular clusters challenges the conventional NS formation theory (Pfahl02). As the vast majority of pulsars detected in globular clusters are MSPs, the crux of the globular cluster NS retention problem is investigating why globular clusters host more MSPs than expected. However, the absence of the MSP $v_{\mathrm{sys}}$ distribution has led to significant uncertainties in the investigation. Furthermore, while the mean 3D space velocities estimated for MSPs by Hobbs05; Gonzalez11 are consistent with our median $v_{\mathrm{sys}}$ (see Section 3), the relatively large astrometric uncertainties in previous works would disproportionately inflate $v_{\perp}$ at the lower end of their $v_{\perp}$ distributions, which is expected to have lowered the predicted indicative rates of MSP retention in globular clusters.
Provided the null hypothesis that MSPs in star clusters follow the same formation channels (and accordingly the same $v_{\mathrm{sys}}$ distribution) as field MSP systems, the $v_{\mathrm{sys}}$ CDF (see Equation 9) derived for field MSP systems can also be seen as the indicative MSP retention fraction as a function of the star cluster central escape velocity $v_{\mathrm{esc}}$. At the nominal globular cluster $v_{\mathrm{esc}}=50$$\mathrm{km\penalty 10000\ s^{-1}}$, we would expect $\approx 14\%$ of MSPs to be retained in the globular cluster, which is no longer inconsistent with the expected retention fraction range of roughly 10% to 20% for binary systems (Drukier96). The indicative MSP retention fraction would further rise to $\gtrsim 48$% for a globular cluster with exceptionally high $v_{\mathrm{esc}}$ of $\gtrsim 100$$\mathrm{km\penalty 10000\ s^{-1}}$ (e.g., Terzan 5). Therefore, our MSP $v_{\mathrm{sys}}$ distribution may be able to account for the large number of MSPs detected in globular clusters, potentially alleviating concerns regarding NS retention in globular clusters. Accordingly, we see no evidence that the aforementioned null hypothesis (that MSPs in star clusters follow the same formation channels as field MSP systems) is violated. More accurate determination of the MSP retention fraction requires careful N-body simulations based on the MSP $v_{\mathrm{sys}}$ distribution, which is beyond the scope of this work.
@@ 6 Summary and Prospects
This work derives the systemic recoil velocity ($v_{\mathrm{sys}}$) distribution of field millisecond pulsar (MSP) systems from the observed $v_{\mathrm{l}}$ (the Galactic-longitude component of the transverse peculiar velocity), assuming that 1) $v_{\mathrm{l}}$ is statistically stable over time (the “stable-$v_{\mathrm{l}}$” assumption), and 2) the $\bm{v}_{\mathrm{sys}}$ directions are uniformly distributed (the “isotropic-$\bm{v}_{\mathrm{sys}}$” assumption). We found that the $v_{\mathrm{sys}}$ distribution can be approximated by a linear combination of three Maxwellian distributions. Using dynamical population synthesis (DPS), we validated the “stable-$v_{\mathrm{l}}$” and evaluated the initial and current Galaxy-wide scale heights of field MSP systems by matching the observed Galactic heights and $v_{\mathrm{l}}$ (the Galactic-latitude component of the transverse peculiar velocity) to the DPS counterparts. The MSP $v_{\mathrm{sys}}$ distribution reported in this work predicts relatively large MSP retention fractions that may account for the large number of MSPs retained in globular clusters.
@@ Acknowledgements.
HD appreciates the helpful comments from the anonymous reviewer, and acknowledges the EACOA Fellowship awarded by the East Asia Core Observatories Association.
@@ Appendix A Selection of the millisecond pulsar sample
We selected the MSP sample from the PSRCAT Catalogue1 (Manchester05) with the help of the psrqpy Python package (Pitkin18). The MSPs were searched using the following criteria:
@@
    $P_{0}<\penalty 10000\$60 ms, where $P_{0}$ represents the spin periods of MSPs, and
@@
    $\dot{P}_{0}<2\times 10^{-17}\,{\mathrm{s\penalty 10000\ s^{-1}}}$, where $\dot{P}_{0}$ denotes the time derivative of $P_{0}$, and
@@
    both proper motion magnitude and parallax are of $3\,\sigma$ significance, with the uncertainty in the proper motion magnitude calculated through the propagation of uncertainties in the right ascension (RA) and the declination components.
@@
There is no clear $P_{0}$ boundary between MSPs and other pulsars. Various $P_{0}$ upper limits up to $\sim 100$ ms have been adopted in different studies of MSPs (e.g. Manchester17; Spiewak22; Ding23). In this work, we adopted the indicative $P_{0}$ upper limit of 60 ms. Additionally, we followed Spiewak22, applying the filter $\dot{P}_{0}<2\times 10^{-17}\,\mathrm{s\penalty 10000\ s^{-1}}$ to exclude recently born ($\lesssim 10$ kyr old) pulsars (such as the Crab pulsar with 34 ms spin period), except for PSR J1417-4402 that does not have a reported $\dot{P}_{0}$. Despite the $\dot{P}_{0}$ being unknown to us, PSR J1417-4402 is suggested to have entered the late recycling phase (Camilo16a), therefore likely being kinematically similar to many other MSPs. Though meeting all the conditions, PSR J1024-0719, which is in an unusually long-period orbit with a main-sequence star, is suggested to be born in (and ejected from) a globular cluster (Kaplan16). Therefore, we removed PSR J1024-0719 from the MSP sample.
@@
As mentioned in Section 2, we look for field MSP systems with $3\,\sigma$ parallax determinations. However, the Gaia-based astrometric results registered in the PSRCAT Catalogue come from Jennings18 based on the Gaia Data Release (DR) 2 (instead of the latest Gaia DR3). Therefore, in practice, we conducted the sample search in three steps. In the first step, we searched for MSPs with a discounted parallax significance threshold of $3\times 0.8=2.4$, where 0.8 is the expected ratio of parallax precision between the Gaia DR3 and the Gaia DR2. In the second step, we updated the astrometric results of Jennings18 to Gaia DR3 utilizing the astroquery package (Ginsburg19). In the final step, we removed MSPs with insignificant ($3\,\sigma$) determinations of parallaxes or proper motions. Compared to parallaxes, proper motions are usually determined at higher significance. The proper motion filter (as part of the criterion iv) only removes one MSP (i.e., PSR J1811-2405) that is too close to the ecliptic plane to acquire precise proper motion in the declination direction with pulsar timing (Ng20). Through the three-step procedure, we reached a sample of 64 field MSP systems, which are compiled in Table 6 along with their most precise proper motion and parallax determinations.
@@
After calculating the transverse peculiar velocities $\bm{v}_{\perp}$ of the 64 MSPs (see Section 2), we found no correlation between $P_{0}$ and $\bm{v}_{\perp}$, hence indicating negligible selection effect of $P_{0}$ on $\bm{v}_{\perp}$.
A.1 High-precision samples
For the selection of the main sample, the thresholds for the significance of proper motions and parallaxes are set to be relatively low (i.e., $3\,\sigma$), as raising the thresholds would inevitably favour high-proper-motion and nearby pulsars. Higher proper motions would likely correspond to larger transverse peculiar velocities $\bm{v}_{\perp}$. On the other hand, given no indication of correlation between $\bm{v}_{\perp}$ and MSP distances, raising the threshold of parallax significance could be rewarding: further lowering astrometric uncertainties, which are dominated by the distance uncertainties, might hold the key to potentially resolving multiple velocity modes. In order to investigate the impact of fractional parallax precision on the velocity analysis, we also adopted higher thresholds of parallax significance for sample selection. For example, with the $5\,\sigma$ and the $7\,\sigma$ thresholds of parallax significances, two “high-precision” samples of 45 and 30 MSPs, respectively, were obtained for comparison with the full sample (of 64 MSPs) (see Table 4).
In the same way as described in Section 2, we compiled $v_{\mathrm{l}}$, $v_{\mathrm{b}}$ and $z$ chains for the high-precision MSP samples, which are displayed in Figure 8 along with the counterparts of the full sample. Compared to the full sample of 64 MSPs, the high-precision MSP samples render consistent $v_{\mathrm{l}}$, $v_{\mathrm{b}}$, and $z$ distributions, while showing increasingly prominent substructures in the $v_{\mathrm{l}}$, $v_{\mathrm{b}}$, and $z$ distributions as sample size drops, likely due to small sample effects.
@@ A.1 High-precision samples
Table 6: Distances $d$, Galactic heights $z$, transverse peculiar velocities $\textit{{v}}_{\perp}\equiv\left(v_{\mathrm{l}},v_{\mathrm{b}}\right)$, and their magnitudes $v_{\perp}$ of the 64 selected millisecond pulsars calculated from their published proper motions ($\mu_{\alpha}$ and $\mu_{\delta}$) and parallaxes $\varpi$, where 1) negative $z$ values correspond to negative Galactic latitudes $b$, and 2) $v_{\mathrm{b}}$ and $v_{\mathrm{l}}$ represent the transverse peculiar velocity components along the directions of $b$ and Galactic longitudes $l$, respectively.